OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

x=5 y=-8

OpenStudy (anonymous):

@anuragtripathi what method did you use

OpenStudy (anonymous):

add the two equations

OpenStudy (anonymous):

after adding you will get x+2x=13+2 3x=15 x=15/3 x=5

OpenStudy (anonymous):

now put the value of x which is 5 in the first equation 5-y=13 shifting y to the right and 13 to the left 5-13=y -8=y

OpenStudy (anonymous):

i hope u get it

OpenStudy (anonymous):

another (simpler) way of doing is to place it in 'y=mx+b' form. graph it then determine where it intersects at. in this case its (5,-8)